The manager of a movie theater found that Saturday's sales were $3675. He knew that a total of 650 tickets were sold Saturday. Adult tickets cost $7.50 and children's tickets cost $4.50. Write equations to represent the number of tickets sold and the amount of money collected. How many of each kind of ticket were sold?
Answer by stanbon(75887)

The manager of a movie theater found that Saturday's sales were $3675. He knew that a total of 650 tickets were sold Saturday. Adult tickets cost $7.50 and children's tickets cost $4.50. Write equations to represent the number of tickets sold and the amount of money collected. How many of each kind of ticket were sold?
---------------------------
Equations:
Quantity Eq:::: a + c = 650
Value Equation:7.5a + 4.5c = 3675
-----------------------------------------
Multiply thru 1st equation by 75:
Multiply thru 2nd equation by 10:
----------------------------
75a + 75c = 75*650
75a + 45c = 36750
---------------------------------
Subtract 2nd from 1st and solve for "c":
30c = 12000
c = 400 (# of children's tickets sold)
-----------------------
Since a+c = 650
a = 250 (# of adult tickets sold)
========================================
Cheers,
Stan H.